OpenStudy (anonymous):

how many moles are present in 34 g of cu(oh)2??? I know it's 0.35 BUT HOW did they get that??? Please help me fast

OpenStudy (anonymous):

HELLO?????? I need HLEP

OpenStudy (cuanchi):

You need to know the molecular mass of the Cu(OH)2. Look in the periodic table the atomic mass for each of the elements in the formula and add them. 1Cu + 2xO + 2 x H = 63.54 + 2 x 16 + 2 x 1 = 97.56 Or you can google "Cu(OH)2" an in Wikipedia you will find the molar mass Now you know that one mole of Cu(OH)2 = 97.56 g 34 g of Cu(OH)2 x 1 mole Cu(OH)2 / 97.56 g Cu(OH)2 = ????? mole of Cu(OH)2